# Continuum Transformers Perform In-Context Learning by Operator Gradient Descent

### v1

@@ 3.1 In-Context Learning Setup
where we now view $\mathcal{M}:\mathcal{V}^{n+1}\rightarrow\mathcal{V}^{n+1}$ as a mask operator acting on the value functions. Note that we present this using a “double-transpose” notation as compared to Equation 3 to follow the mathematical convention of writing an operator to the left of the function it is acting upon. Notably, in the setting of in-context learning for PDEs, the pairs $(f^{(j)},u^{(j)})$ are generally time rollout pairs $(f^{(i)},u^{(i)})=(u^{((i-1)\Delta t)},u^{(i\Delta t)})$ for some time increment $\Delta t$ as discussed in Section 2.3. We, thus, assume $f^{(i)},u^{(i)}\in\mathcal{X}$, that is, that they lie in a single Hilbert space. Following the formalization established in [37], we then seek to learn in-context a map $\mathcal{X}\rightarrow\mathcal{X}$, for which we construct a context window consisting of training pairs $z^{(i)}\in\mathcal{Z}=\mathcal{X}\oplus\mathcal{X}$ as
@@
where $\{(f^{(i)},u^{(i)})\}_{i=1}^{n}$ are $n$ input-output pairs, $z^{(i)}:=(f^{(i)},u^{(i)})$, and $\mathcal{T}:\mathcal{Z}^{n+1}\rightarrow\mathcal{Z}^{n+1}$. We predict $u^{(n+1)}$ as $-[\mathcal{T}(z)]_{2,n+1}$. As in the typical ICL formalization, the key, query, and nonlinearity operators only act upon the input functions. That is, denoting $X_{0}=[f^{(1)},f^{(2)},...,f^{(n)},f^{(n+1)}]\in\mathcal{X}^{n+1}$ as the vector of input functions and $X_{\ell}\in\mathcal{X}^{n+1}$ as the first row of the matrix $Z_{\ell}$,
$Z_{\ell+1}=Z_{\ell}+\left(\widetilde{H}(\mathcal{W}_{q,\ell}X_{\ell},\mathcal{W}_{k,\ell}X_{\ell})\mathcal{M}(\mathcal{W}_{v,\ell}Z_{\ell})^{T}\right)^{T}.$ (8)
Notably, the layer $\ell$ query and key block operators $\mathcal{W}_{q,\ell}:\mathcal{X}^{n+1}\to\mathcal{Q}^{n+1}$ and $\mathcal{W}_{k,\ell}:\mathcal{X}^{n+1}\to\mathcal{K}^{n+1}$ only act on the restricted input space, whereas $\mathcal{W}_{v,\ell}:\mathcal{Z}^{n+1}\to\mathcal{Z}^{n+1}$ acts on the full space. The mask block operator matrix $\mathcal{M}$ is set as $\begin{bmatrix}I_{n\times n}&0\\ 0&0\end{bmatrix}$, with $I_{n\times n}$ denoting the $n\times n$ block identity operator. We denote the negated in-context prediction for an input $f^{(n+1)}:=f$ after $\ell$ layers by
@@ 3.2 In-Context Learning Occurs via Gradient Descent
We now show that continuum transformers can perform in-context learning by implementing operator gradient descent. For in-context samples $\{(f^{(i)},u^{(i)})\}^{n}_{i=1}$, the in-context task is to find
$O^{*}:=\min_{O\in\mathcal{O}}L(O):=\min_{O\in\mathcal{O}}\sum_{i=1}^{n}\|u^{(i)}-Of^{(i)}\|^{2}_{\mathcal{X}}.$ (11)
Naturally, such an operator learning problem can be solved by iteratively updating an estimate $\widehat{O}_{t+1}=\widehat{O}_{t}-\eta\nabla_{O}L(O)$. Our result, therefore, roughly states that the prediction of $u^{(n+1)}$ by the $\ell$-th layer of a continuum transformer, with particular choices of key, query, and value parameters, is exactly $\widehat{O}_{\ell}f$ for any test function $f$. We defer relevant definitions and the full proof of the theorem to Appendix C. We, however, here highlight novelties of the proof strategy in demonstrating this result. The strategy involves deriving an explicit form of the operator gradient descent steps and then showing, under a specific choice of $(\mathcal{W}_{v},\mathcal{W}_{q},\mathcal{W}_{k})$, inference through a layer of the continuum transformer recovers this explicitly computed operator gradient expression (see Appendix B). While the general strategy parallels that of [27], the explicit construction of the gradient descent expression over operator spaces is not as directly evident as it is over finite-dimensional vector spaces. In particular, we had to invoke a generalized form of the Representer Theorem (as provided by [38]) to obtain this result, as compared to its classical form leveraged in [27].
@@
Let $\kappa:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{L}(\mathcal{X})$ be an arbitrary operator-valued kernel and $\mathcal{O}$ be the operator RKHS induced by $\kappa$. Let $\{(f^{(i)},u^{(i)})\}^{n}_{i=1}$ and $L$ be as defined by Equation 11. Let $O_{0}=\mathbf{0}$ and let $O_{\ell}$ denote the operator obtained from the $\ell$-th operator-valued gradient descent sequence of $L$ with respect to $\|\cdot\|_{\mathcal{O}}$ as defined in Equation 25. Then there exist scalar step sizes $r_{0}^{\prime},\ldots,r_{m}^{\prime}$ such that if, for an $m$-layer continuum transformer $\mathcal{T}:=\mathcal{T}_{m}\circ...\circ\mathcal{T}_{0}$, $[\widetilde{H}(U,W)]_{i,j}=\kappa(u^{(i)},w^{(j)})$, $\mathcal{W}_{v,\ell}=\begin{pmatrix}0&0\\ 0&-r^{\prime}_{\ell}I\end{pmatrix}$, $\mathcal{W}_{q,\ell}=I$, and $\mathcal{W}_{k,\ell}=I$ for each $\ell=0,...,m$, then for any $f\in\mathcal{X}$,
@@ 3.3 In-Context Learning Recovers the Best Linear Unbiased Predictor
In the finite-dimensional setting, it was shown in [27] that if $\{y^{(j)}\}$ are outputs from the marginal of a Gaussian process with kernel $\kappa(\cdot,\cdot)$ for inputs $\{x^{(j)}\}$, a transformer of depth $m\rightarrow\infty$ with $[H(U,W)]_{i,j}=\kappa(u^{(i)},w^{(j)})$ recovers the Bayes optimal predictor in-context for a window of $\{(x^{(j)},y^{(j)})\}$. Roughly speaking, this result demonstrates the optimality of recovering the true $f$ in-context if the distribution on the function space from which $f$ was sampled is Gaussian. We, therefore, seek to establish a similar result for the setting of operator-valued kernels, which requires defining a Gaussian measure on the operator space from which the true $O$ is sampled.
Notably, paralleling how the finite marginals of finite-dimensional Gaussian processes induce distributions over functions, characterizing the finite marginals of a Hilbert space-valued Gaussian process induces a distribution over operators. Doing so, however, requires the formalization of a Gaussian measure on Hilbert spaces. A Hilbert space-valued random variable $F$ is said to be Gaussian if $\langle\varphi,F\rangle_{\mathcal{X}}$ is Gaussian for any $\varphi\in\mathcal{X}$ [39]. We now present the generalization of Gaussian processes to Hilbert spaces from [40], adjusting the notation to match that used herein.
@@
$\kappa$ Gaussian Random Variable in Hilbert Space (Definition 4.1 of [40]) Given an operator-valued kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{X})$, we say $U|F\sim\mathcal{N}(\mathbf{0},\mathbf{K}(F)),$ where $U=[u^{(1)},\ldots,u^{(n+1)}],F=[f^{(1)},\ldots,f^{(n+1)}]$ and $[\mathbf{K}(F)]_{i,j}:=\kappa(f^{(i)},f^{(j)})$ if, for all $v^{(1)},v^{(2)}\in\mathcal{X}$ and $(f^{(i)},u^{(i)}),(f^{(j)},u^{(j)})\in\mathcal{X}^{2}$,
@@
$\langle v^{(1)},u^{(i)}\rangle_{\mathcal{X}}\sim\mathcal{N}\left(0,\langle v^{(1)},\kappa(f^{(i)},f^{(i)})v\rangle_{\mathcal{X}}\right).$ (13)
Such Hilbert-space valued GPs have been leveraged in Hilbert-space kriging, notably studied in [39, 41]. It was there of interest to characterize the resulting GP estimator. In turn, they proved that the Best Linear Unbiased Predictor (BLUP) precisely coincides with the Bayes optimal predictor in the MSE sense, meaning the $f:\mathcal{X}^{n}\to\mathcal{X}$ amongst all measurable functions that minimizes $\mathbb{E}[\|X_{n+1}-f(\mathbf{X})\|^{2}_{\mathcal{X}}]$, for zero-mean jointly Gaussian random variables $\mathbf{X}=(X_{1},\ldots,X_{n})\in\mathcal{X}^{n}$ and $X_{n+1}\in\mathcal{X}$, is the BLUP. Their formal statement is provided for reference in Section D.1.
We now show that the in-context learned operator is the Best Linear Unbiased Predictor as $m\rightarrow\infty$ and thus, by Theorem D.1, the Bayes optimal predictor if the observed data are $\kappa$ Gaussian random variables and if $\tilde{H}$ matches $\kappa$. We again highlight here the technical approach and main challenges in establishing this result and defer the full presentation to Section D.2. Roughly speaking, the approach relied on establishing that the infinite composition $\mathcal{T}_{\infty}:=...\circ\mathcal{T}_{m}\circ...\circ\mathcal{T}_{0}$ converges to a well-defined operator and that this matches the form of the BLUP known from Hilbert space kriging literature. The need to carefully handle this compositional convergence and connection to Hilbert space kriging are novel technical contributions of this result.
@@
Let $F=[f^{(1)},\ldots,f^{(n+1)}],U=[u^{(1)},\ldots,u^{(n+1)}]$. Let $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{X})$ be an operator-valued kernel. Assume that $U|F$ is a $\kappa$ Gaussian random variable per Definition 3.2. Let the activation function $\tilde{H}$ of the attention layer be defined as $[\tilde{H}(U,W)]_{i,j}:=\kappa(u^{(i)},w^{(j)})$. Consider the operator gradient descent in Theorem 3.1. Then as the number of layers $m\to\infty$, the continuum transformer’s prediction at layer $m$ approaches the Best Linear Unbiased Predictor that minimizes the in-context loss in Equation 10.
3.4 Pre-Training Continuum Transformers Can Converge to Gradient Descent Parameters
@@
We now demonstrate that the aforementioned parameters that result in operator gradient descent from Theorem 3.1 are in fact minimizers of the continuum transformer training objective. This, in turn, suggests that under such training, the continuum transformer parameters will converge to those exhibited in the previous section. Similar results have been proven for the standard Transformer architecture, such as in [27]. Demonstrating our results, however, involved highly non-trivial changes to their proof strategy. In particular, defining gradient flows over the space of functionals of a Hilbert space requires the notion of Frechet-differentiability, which is more general than taking derivatives with respect to a matrix. Additionally, rewriting the training objective with an equivalent expectation expression (see Appendix E) required careful manipulation of the covariance operators of the data distributions as briefly described in the symmetry discussion that follows.
@@
The formal statement of this result is given in Theorem 3.6. This result characterizes a stationary point of the optimization problem in Equation 14 when the value operators $W_{v,\ell}$ have the form $\begin{bmatrix}0&0\\ 0&r_{\ell}\end{bmatrix}$ and the $W_{q,\ell}$ and $W_{k,\ell}$ operators have a form relating to the symmetry of the data distribution, as fully described below. In particular, if the symmetry is characterized by a self-adjoint invertible operator $\Sigma$, we establish that there exists a fixed point of the form $W_{q,\ell}=b_{\ell}\Sigma^{-1/2}$ and $W_{k,\ell}=c_{\ell}\Sigma^{-1/2}$ for some constants $\{b_{\ell}\}$ and $\{c_{\ell}\}$. Notably, this fixed point relates to Section 3.2 and Section 3.3, since in the particular case where $\Sigma=I$, we recover the parameter configuration under which the transformer performs functional gradient descent.
@@
As alluded to, this proof relies on some technical assumptions on the $F$ and $U|F$ distributions and transformer nonlinearity; such assumptions are typical of such optimization analyses, as seen in related works such as [27, 42, 43]. As mentioned, the proof proceeds by studying gradient flow dynamics of the $\mathcal{W}_{k,\ell}$ and $\mathcal{W}_{q,\ell}$ operators over the $\mathcal{P}(F,U)$ distribution. Direct analysis of such gradient flows, however, becomes analytically unwieldy under unstructured distributions $\mathcal{P}(F,U)$. We, therefore, perform the analysis in a frame of reference rotated by $\Sigma^{-1/2}$. By E.3, this rotated frame preserves the $\mathcal{P}(F,U)$ distribution and by E.4 also the attention weights computed by the continuum transformer, allowing us to make conclusions on the original setting after performing the analysis in this rotated coordinate frame. We demonstrate in Section 4.3 that kernels that are often of interest satisfy E.4. We defer the full proof to Appendix E.
@@
Suppose 3.4 and 3.5 hold. Let $f(r,\mathcal{W}_{q},\mathcal{W}_{k}):=L\left(\mathcal{W}_{v,\ell}=\left\{\begin{bmatrix}0&0\\ 0&r_{\ell}\end{bmatrix}\right\}_{\ell=0,\ldots,m},\mathcal{W}_{q,\ell},\mathcal{W}_{k,\ell}\right)$, where $L$ is as defined in Equation 10. Let $\mathcal{S}\subset\mathcal{O}^{m+1}\times\mathcal{O}^{m+1}$ denote the set of $(\mathcal{W}_{q},\mathcal{W}_{k})$ operators with the property that $(\mathcal{W}_{q},\mathcal{W}_{v})\in\mathcal{S}$ if and only if for all $\ell\in\{0,\ldots,m\}$, there exist scalars $b_{\ell},c_{\ell}\in\mathbb{R}$ such that $\mathcal{W}_{q,\ell}=b_{\ell}\Sigma^{-1/2}$ and $\mathcal{W}_{k,\ell}=c_{\ell}\Sigma^{-1/2}$. Then

### v2

@@ 3.1 In-Context Learning Model
where we now view $\mathcal{M}:\mathcal{V}^{n+1}\rightarrow\mathcal{V}^{n+1}$ as a mask operator acting on the value functions. Note that we present this using a “double-transpose” notation as compared to Equation 2 to follow the mathematical convention of writing an operator to the left of the function it is acting upon. Notably, in the setting of in-context learning for PDEs, the pairs $(f^{(j)},u^{(j)})$ are generally time rollout pairs $(f^{(i)},u^{(i)})=(u^{((i-1)\Delta t)},u^{(i\Delta t)})$ for some time increment $\Delta t$, elaborated upon more in Appendix B. We, thus, assume $f^{(i)},u^{(i)}\in\mathcal{X}$, that is, that they lie in a single Hilbert space. Following the formalization established in [34], we then seek to learn in-context a map $\mathcal{X}\rightarrow\mathcal{X}$, for which we construct a context window consisting of training pairs $z^{(i)}\in\mathcal{Z}=\mathcal{X}\oplus\mathcal{X}$:
@@
where $\{(f^{(i)},u^{(i)})\}_{i=1}^{n}$ are $n$ input-output pairs, $z^{(i)}:=(f^{(i)},u^{(i)})$, and $\mathcal{T}:\mathcal{Z}^{n+1}\rightarrow\mathcal{Z}^{n+1}$. We predict $u^{(n+1)}$ as $\mathcal{T}(z)]_{2,n+1}$. As in the typical ICL formalization, the key, query, and nonlinearity operators only act upon the input functions. That is, denoting $X_{0}=[f^{(1)},f^{(2)},...,f^{(n)},f^{(n+1)}]\in\mathcal{X}^{n+1}$ as the vector of input functions and $X_{\ell}\in\mathcal{X}^{n+1}$ similarly as the first row of the matrix $Z_{\ell}$,
$Z_{\ell+1}=Z_{\ell}+\left(\widetilde{H}(\mathcal{W}_{q,\ell}X_{\ell},\mathcal{W}_{k,\ell}X_{\ell})\mathcal{M}(\mathcal{W}_{v,\ell}Z_{\ell})^{T}\right)^{T},$ (7)
with the layer $\ell$ query and key block operators $\mathcal{W}_{q,\ell}:\mathcal{X}^{n+1}\to\mathcal{Q}^{n+1}$ and $\mathcal{W}_{k,\ell}:\mathcal{X}^{n+1}\to\mathcal{K}^{n+1}$ only acting on the restricted input space and $\mathcal{W}_{v,\ell}:\mathcal{Z}^{n+1}\to\mathcal{Z}^{n+1}$ on the full space. The mask block operator matrix $\mathcal{M}$ is set as $\begin{bmatrix}I_{n\times n}&0\\ 0&0\end{bmatrix}$, with $I_{n\times n}$ denoting the $n\times n$ block identity operator. We denote the negated in-context prediction for an input $f^{(n+1)}:=f$ after $\ell$ layers by
@@
Notably, the very modeling of the generalized continuum layer in Equation 5 by introducing $H:\mathcal{Q}^{n+1}\times\mathcal{K}^{n+1}\to(\mathcal{L}(\mathcal{V}))^{(n+1)\times(n+1)}$ was a non-obvious generalization of Equation 3 in introducing this operator-valued nonlinearity. The more natural thing would have been to simply consider a scalar-valued function “similarity” measure, i.e. defining $H:\mathcal{Q}^{n+1}\times\mathcal{K}^{n+1}\to\mathbb{R}^{(n+1)\times(n+1)}$. However, critically, doing so does not lend itself to characterization in an RKHS, upon which our analysis over the remaining section is based. Thus, the mathematical framing of this problem as above is a worthwhile contribution that the broader community can use for further analysis.
@@ 3.2 In-Context Learning Occurs via Gradient Descent
We now show that continuum transformers can perform in-context learning by implementing operator gradient descent. For in-context samples $\{(f^{(i)},u^{(i)})\}^{n}_{i=1}$, the in-context task is to find $O^{*}:=\min_{O\in\mathcal{O}}L(O):=\min_{O\in\mathcal{O}}\sum_{i=1}^{n}\|u^{(i)}-Of^{(i)}\|^{2}_{\mathcal{X}}$. Naturally, such an operator learning problem can be solved by iteratively updating an estimate ${O}_{\ell+1}={O}_{\ell}-\eta_{\ell}\nabla L({O}_{\ell})$. Our result, therefore, roughly states that the prediction of $u^{(n+1)}$ by the $\ell$-th layer of a continuum transformer, with particular choices of key, query, and value parameters, is exactly ${O}_{\ell}f$ for any test function $f$. We defer the full proof of the theorem to Appendix E but here highlight novelties of the proof strategy in demonstrating this result. The strategy involves deriving an explicit form of the operator gradient descent steps and then showing, under a specific choice of $(\mathcal{W}_{v},\mathcal{W}_{q},\mathcal{W}_{k})$, inference through a layer of the continuum transformer recovers this explicitly computed operator gradient expression (see Appendix D). While the general strategy parallels that of [24], the explicit construction of the gradient descent expression over operator spaces is not as directly evident as it is over finite-dimensional vector spaces. In particular, we had to invoke a generalized form of the Representer Theorem to obtain this result, as compared to its classical form leveraged therein.
@@
Let $\kappa:\mathcal{X}\times\mathcal{X}\rightarrow\mathcal{L}(\mathcal{X})$ be an arbitrary operator-valued kernel and $\mathcal{O}$ be the operator RKHS induced by $\kappa$. Let $\{(f^{(i)},u^{(i)})\}^{n}_{i=1}$ and $L(O):=\sum_{i=1}^{n}\|u^{(i)}-Of^{(i)}\|^{2}_{\mathcal{X}}$. Let $O_{0}=\mathbf{0}$ and let $O_{\ell}$ denote the operator obtained from the $\ell$-th operator-valued gradient descent sequence of $L$ with respect to $\|\cdot\|_{\mathcal{O}}$ as defined in Equation 22. Then there exist scalar step sizes $r_{0}^{\prime},\ldots,r_{m}^{\prime}$ such that if, for an $m$-layer continuum transformer $\mathcal{T}:=\mathcal{T}_{m}\circ...\circ\mathcal{T}_{0}$, $\widetilde{H}(U,W)]_{i,j}=\kappa(u^{(i)},w^{(j)})$, $\mathcal{W}_{v,\ell}=\begin{pmatrix}0&0\\ 0&-r^{\prime}_{\ell}I\end{pmatrix}$, $\mathcal{W}_{q,\ell}=I$, and $\mathcal{W}_{k,\ell}=I$ for each $\ell=0,...,m$, then for any $f\in\mathcal{X}$,
@@ 3.3 In-Context Learning Recovers the Best Linear Unbiased Predictor
In the finite-dimensional setting, it was shown in [24] that if $\{y^{(j)}\}$ are outputs from the marginal of a Gaussian process with kernel $\kappa(\cdot,\cdot)$ for inputs $\{x^{(j)}\}$, a transformer of depth $m\rightarrow\infty$ with $H(U,W)]_{i,j}=\kappa(u^{(i)},w^{(j)})$ recovers the Bayes optimal predictor in-context for a window of $\{(x^{(j)},y^{(j)})\}$. Roughly speaking, this result demonstrates the optimality of recovering the true $f$ in-context if the distribution on the function space from which $f$ was sampled is Gaussian. We, therefore, seek to establish a similar result for the setting of operator-valued kernels, which requires defining a Gaussian measure on the operator space from which the true $O$ is sampled.
Notably, paralleling how the finite marginals of finite-dimensional Gaussian processes induce distributions over functions, characterizing the finite marginals of a Hilbert space-valued Gaussian process induces a distribution over operators. Doing so, however, requires the formalization of a Gaussian measure on Hilbert spaces. A Hilbert space-valued random variable $F$ is said to be Gaussian if $\langle\varphi,F\rangle_{\mathcal{X}}$ is Gaussian for any $\varphi\in\mathcal{X}$ [35]. We now present the generalization of Gaussian processes to Hilbert spaces from [36].
@@
$\kappa$ Gaussian Random Variable in Hilbert Space (Definition 4.1 of [36]) Given an operator-valued kernel $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{X})$, we say $U|F\sim\mathcal{N}(\mathbf{0},\mathbf{K}(F)),$ where $U=[u^{(1)},\ldots,u^{(n+1)}],F=[f^{(1)},\ldots,f^{(n+1)}]$ and $\mathbf{K}(F)]_{i,j}:=\kappa(f^{(i)},f^{(j)})$ if, for all $v^{(1)},v^{(2)}\in\mathcal{X}$ and $(f^{(i)},u^{(i)}),(f^{(j)},u^{(j)})\in\mathcal{X}^{2}$,
@@
$\langle v^{(1)},u^{(i)}\rangle_{\mathcal{X}}\sim\mathcal{N}\left(0,\langle v^{(1)},\kappa(f^{(i)},f^{(i)})v^{(1)}\rangle_{\mathcal{X}}\right).$ (11)
Such Hilbert-space valued GPs have been leveraged in Hilbert-space kriging, notably studied in [35, 37]. They proved that the Best Linear Unbiased Predictor (BLUP) precisely coincides with the Bayes optimal predictor in the MSE sense, meaning the $f:\mathcal{X}^{n}\to\mathcal{X}$ amongst all measurable functions that minimizes $\mathbb{E}[\|X_{n+1}-f(\mathbf{X})\|^{2}_{\mathcal{X}}]$, for zero-mean jointly Gaussian random variables $\mathbf{X}=(X_{1},\ldots,X_{n})\in\mathcal{X}^{n}$ and $X_{n+1}\in\mathcal{X}$, is the BLUP. Their formal statement is provided for reference in Section F.1.
We now show that the in-context learned operator is the Best Linear Unbiased Predictor as $m\rightarrow\infty$ and thus, by Theorem F.1, Bayes optimal if the observed data are $\kappa$ Gaussian random variables and if $\widetilde{H}$ matches $\kappa$. We again highlight here the approach and main innovations in establishing this result and defer the full presentation to Section F.2. Roughly, the approach relied on establishing that the infinite composition $\mathcal{T}_{\infty}:=...\circ\mathcal{T}_{m}\circ...\circ\mathcal{T}_{0}$ converges to a well-defined operator and that this matches the BLUP known from Hilbert space kriging literature. The careful handling of this compositional convergence and connection to Hilbert space kriging are novel technical contributions.
@@
Let $F=[f^{(1)},\ldots,f^{(n+1)}],U=[u^{(1)},\ldots,u^{(n+1)}]$. Let $\kappa:\mathcal{X}\times\mathcal{X}\to\mathcal{L}(\mathcal{X})$ be an operator-valued kernel. Assume that $U|F$ is a $\kappa$ Gaussian random variable per Definition 3.2. Let the activation function $\widetilde{H}$ of the attention layer be defined as $\widetilde{H}(U,W)]_{i,j}:=\kappa(u^{(i)},w^{(j)})$. Consider the operator gradient descent in Theorem 3.1. Then as the number of layers $m\to\infty$, the continuum transformer’s prediction at layer $m$ approaches the Best Linear Unbiased Predictor that minimizes the in-context loss in Equation 9.
3.4 Pre-Training Can Converge to Gradient Descent Parameters
@@
We now demonstrate that the aforementioned parameters that result in operator gradient descent from Theorem 3.1 are in fact minimizers of the continuum transformer training objective. This, in turn, suggests that under such training, the continuum transformer parameters will converge to those exhibited in the previous section. Similar results have been proven for the standard Transformer architecture, such as in [24]. Demonstrating our results, however, involved highly non-trivial changes to their proof strategy. In particular, defining gradient flows over the space of functionals of a Hilbert space requires the notion of Frechet-differentiability, which is more general than taking derivatives with respect to a matrix. Additionally, rewriting the training objective with an equivalent expectation expression (see Appendix G) required careful manipulation of the covariance operators of the data distributions as described in the symmetry discussion that follows.
@@
The formal statement of this result is given in Theorem 3.6. This result characterizes a stationary point of the optimization problem in Equation 12 when the value operators $W_{v,\ell}$ have the form $\begin{bmatrix}0&0\\ 0&r_{\ell}I\end{bmatrix}$ and the $W_{q,\ell}$ and $W_{k,\ell}$ operators have a form relating to the symmetry of the data distribution, as fully described below. If the symmetry is characterized by a self-adjoint invertible operator $\Sigma$, we establish that there exists a fixed point of the form $W_{q,\ell}=b_{\ell}\Sigma^{-1/2}$ and $W_{k,\ell}=c_{\ell}\Sigma^{-1/2}$ for some constants $\{b_{\ell}\}$ and $\{c_{\ell}\}$. This fixed point relates to Section 3.2 and Section 3.3, since if $\Sigma=I$, we recover the parameter configuration under which functional gradient descent is performed.
@@
This proof relies on some technical assumptions on the $F$ and $U|F$ distributions and transformer nonlinearity; such assumptions are typical of such optimization analyses, as seen in related works such as [24, 38, 39]. As mentioned, the proof proceeds by studying gradient flow dynamics of the $\mathcal{W}_{k,\ell}$ and $\mathcal{W}_{q,\ell}$ operators over the $\mathcal{P}(F,U)$ distribution. Direct analysis of such gradient flows, however, becomes analytically unwieldy under unstructured distributions $\mathcal{P}(F,U)$. We, therefore, perform the analysis in a frame of reference rotated by $\Sigma^{-1/2}$. By G.3, this rotated frame preserves the $\mathcal{P}(F,U)$ distribution and by G.4 also the attention weights computed by the continuum transformer, allowing us to make conclusions on the original setting after performing the analysis in this rotated coordinate frame. We demonstrate in Section 5.3 that common kernels satisfy G.4. We defer the full proof to Appendix G.
@@
(Rotational symmetry) Let $P_{F}$ be the distribution of $F=[f^{(1)},...,f^{(n+1)}]$ and $\mathbb{K}(F)=\mathbb{E}_{U|F}[U\otimes U]$. We assume that there exists a self-adjoint, invertible operator $\Sigma:\mathcal{X}\to\mathcal{X}$ such that for any unitary operator $\mathcal{M}$, $\Sigma^{1/2}\mathcal{M}\Sigma^{-1/2}F\stackrel{{\scriptstyle d}}{{=}}F$ and $\mathbb{K}(F)=\mathbb{K}(\Sigma^{1/2}\mathcal{M}\Sigma^{-1/2}F)$.
@@
For any $F_{1},F_{2}\in\mathcal{X}^{n+1}$ and any operator $S:\mathcal{X}\to\mathcal{X}$ with an inverse $S^{-1}$, the function $\widetilde{H}$ satisfies $\widetilde{H}(F_{1},F_{2})=\widetilde{H}(S^{*}F_{1},S^{-1}F_{2})$.
@@
Suppose 3.4 and 3.5 hold. Let $f(r,\mathcal{W}_{q},\mathcal{W}_{k}):=L\left(\mathcal{W}_{v,\ell}=\left\{\begin{bmatrix}0&0\\ 0&r_{\ell}I\end{bmatrix}\right\}_{\ell=0,\ldots,m},\mathcal{W}_{q,\ell},\mathcal{W}_{k,\ell}\right)$, where $L$ is as defined in Equation 9. Let $\mathcal{S}\subset\mathcal{O}^{m+1}\times\mathcal{O}^{m+1}$ denote the set of $(\mathcal{W}_{q},\mathcal{W}_{k})$ operators with the property that $(\mathcal{W}_{q},\mathcal{W}_{v})\in\mathcal{S}$ if and only if for all $\ell\in\{0,\ldots,m\}$, there exist scalars $b_{\ell},c_{\ell}\in\mathbb{R}$ such that $\mathcal{W}_{q,\ell}=b_{\ell}\Sigma^{-1/2}$ and $\mathcal{W}_{k,\ell}=c_{\ell}\Sigma^{-1/2}$. Then
@@
4 Related Works
We were interested herein in providing a theoretical characterization of the in-context learning exhibited by continuum attention-based ICONs, paralleling that done for finite-dimensional transformers [3, 4, 5]. While previous works have yet to formally characterize ICL for continuum transformers, a recent work [34] began a line of inquiry in this direction. Their work, however, studied a fundamentally distinct aspect of functional ICL, characterizing the sample complexity and resulting generalization for linear elliptic PDEs.
Most relevant in the line of works characterizing finite-dimensional ICL is [24]. Loosely speaking, they demonstrated that, if a kernel $\kappa(\cdot,\cdot)$ is defined with $\mathbb{H}$ denoting its associated RKHS and a transformer is then defined with a specific $W_{k},W_{q},W_{v}$ and $H(Q,K)]_{i,j}=\kappa(Q_{i},K_{j})$ from Equation 2, an inference pass through $m$ layers of such a transformer predicts $y^{(n+1)}_{i}$ from Equation 4 equivalently to $\widehat{f}(x^{(n+1)}_{i})$, where $\widehat{f}$ is the model resulting from $m$ steps of functional gradient descent in $\mathbb{H}$. They then demonstrated that the predictor learned by such ICL gradient descent is Bayes optimal under particular circumstances. These results are provided in Appendix C.
Notably, these results required non-trivial changes to be generalized to operator RKHSs as we did herein; as discussed, such mathematical tooling is more general than its application to the setting considered herein. In particular, our approach to performing the analysis in the infinite-dimensional function space directly without requiring discretization required formalizing several notions to rigorously justify being able to lift the proofs from finite to infinite dimensions. We view this as a significant contribution to the community. Now that we have gone through this formalism, the broader community can directly use these analysis strategies to further study in-context learning or unrelated inquiries. This is highly valuable, since it suggests that theoreticians can often follow their finite-dimensional intuitions and defer to our infinite-dimensional results to rigorously justify their results. In other words, our proof strategies suggest to other researchers working on similar problems that they need not be bogged down in the details of the error convergence minutiae that often arise in approaches relying on finite projections and can instead work with these clean abstractions.
Similarly, this work opens up the space for who can contribute to further the theoretical study of operator ICL. Much of the classical optimization community, for instance, may not be intimately familiar with the mathematical formalisms required around operator spaces. However, with our formalism, they can provide insights with little change from how they would approach finite-dimensional analyses. We, therefore, believe this framework, consisting of the generalized transformer and characterization of its optimization, is a worthwhile contribution to the theory community.